can you please help... tomorrow is my math exam and I cant understand this sum!!
2. At present the sum of Monica’s age and her daughter’s age is 44 years. After
2 years, Monika’s age will be 3 times that of her daughter’s age. Find their
present ages.

Answer by macston(5194)

M=Monika's age; D=daughter's age
M+D=44 years
M=44yrs-D
M+2yrs=3(D+2yrs) Substitute for M
44yrs-D+2yrs=3D+6yrs Subtract 6 yrs from each side.
40 yrs-D=3D Add D to each side.
40 yrs=4D Divide each side by 4.
10 yrs=D ANSWER 1: The daughter is 10 years old.
M=44 yrs-D=44yrs-10yrs=34 years ANSWER 2: Monika is 34 years old.
CHECK:
M+2 yrs=3(D+2yrs)
34yrs+2yrs=3(10yrs+2yrs)
36yrs=3(12yrs)
36yrs=36yrs